# Effects of Mosaic Crystal Instrument Functions on X-ray Thomson Scattering Diagnostics

### v1

@@ Abstract
Mosaic crystals, with their high integrated reflectivities, are widely-employed in spectrometers used to diagnose high energy density systems. X-ray Thomson scattering (XRTS) has emerged as a powerful diagnostic tool of these systems, providing in principle direct access to important properties such as the temperature via detailed balance. However, the measured XRTS spectrum is broadened by the spectrometer instrument function (IF), and without careful consideration of the IF one risks misdiagnosing system conditions. Here, we consider in detail the IF of mosaic crystals and how the broadening varies across the spectrometer. Notably, we find a strong asymmetry in the shape of the IF towards higher energies. As an example, we consider the effect on the inferred temperature, and find that it can be overestimated if an approximate symmetric IF is used. We therefore expect a detailed consideration of the full IF will have an important impact on system properties inferred via XRTS in both forward modelling and model-free approaches.
@@ I Introduction
where ${\bm{q}}$ is the scattering vector, $E$ is the energy scattered photon measured on the spectrometer, $E_{0}$ is the energy of the incident photon, and $\hbar\omega=E_{0}-E$ is the energy loss. The SIF contains all the effects that broaden the measured XRTS spectrum, specifically the spectrum of the incident probe beam (the source) and the instrument function of the spectrometer. The temperature $k_{B}T=\beta^{-1}$ of a system in thermal equilibrium, for example, may in principle be extracted using the detailed balance relation of the DSF:
@@
An x-ray spectrometer typically consists of two components: a dispersive crystal and a photosensitive optic. As the light emitted from a target is polychromatic, in order to measure the spectrum the different photon energies need to be separated from each other. This is often done using a crystal with an appropriate lattice spacing to disperse the x-rays in space, which are then detected by the photosensitive optic such as a charge-coupled device (CCD) camera Gruner, Tate, and Eikenberry (2002). The spatial position of a photon on the camera can be related to its energy via Bragg’s law and the geometry of the spectrometer v. Hámos (1932); Zachariasen (1967).
Mosaic crystals are a popular choice for the dispersive crystal. While perfect crystals consist of uniform layers of atoms, mosaic crystals consist of small perfect crystallites with normal vectors randomly distributed about the surface normal Zachariasen (1967); Freund (1988); Legall et al. (2006a); Gerlach et al. (2015). As a result, in contrast to perfect crystals where the Bragg condition will only be satisfied along a single plane through the crystal, on a mosaic crystal it can be satisfied anywhere. The solid angle coverage of a mosaic crystal is therefore substantially higher than that of a perfect crystal, resulting in a much higher integrated reflectivity Legall et al. (2006b, 2009). In the diagnostics of WDM systems in the laboratory, where these systems are often transient, this higher reflectivity has led to the widespread adoption of mosaic crystals as they collect more photons in a given shot. But this increased reflectivity comes at the cost of spectral resolution as the distribution of the crystallites also broadens the measured spectrum. For example, in widely-used Highly Annealed Pyrolytic Graphite (HAPG) crystals, the distribution of crystallites is approximately Lorentzian Zastrau et al. (2013); Gerlach et al. (2015), and therefore the crystal broadening of spectra has generally been treated as a convolution of the emitted spectrum with a Lorentzian or a Voigt profile Preston et al. (2020); MacDonald et al. (2021).
However, recent ray tracing simulations indicate that the actual broadening by mosaic crystals is very asymmetric towards higher energies Šmíd, Pan, and Falk (2021). An asymmetric SIF would have consequences on conditions derived from XRTS measurements if an approximate SIF that is symmetric – or one without sufficient asymmetry – were used in its place. A clear example would be that the temperature extracted via detailed balance would be incorrect as the upshifted side of the DSF would be inferred to be larger to compensate for the increase in intensity by the SIF.
@@
In this paper, we present a systematic analysis of the broadening by mosaic crystals and the effect it has on plasma conditions inferred in XRTS measurements. To do so, we use Schlesiger et al.’s multi-reflection mosaic crystal model Schlesiger et al. (2017) to calculate the differential reflectivity along paths a photon can take from the source to the camera. The model is found to have good agreement with experimental measurements taken at the European XFEL on ambient Al and polypropylene. Notably, we find the mosaic broadening and depth broadening result in highly asymmetric broadening towards the higher energies. The instrument function is also found to be extremely extended, with substantial intensity even more than 100 eV away from the central photon energy. Furthermore, it is demonstrated that such strong asymmetry in the beam has implications on conditions inferred from XRTS via forward modelling and the model-free method if an approximate symmetric SIF is used. Finally, we investigate the validity of the approximation of the SIF as a convolution problem rather than a kernel problem.
@@ II Broadening by Mosaic Crystals
Figure 1: A two-dimensional schematic of a von Hámos detector with a mosaic crystal. In full von Hámos geometry, the crystal is cylindrically bent with a radius of curvature $r$, while $X_{\rm cam}$ is the distance from the source to a point on the camera. Mosaic focusing in this geometry is represented by photons incident on the crystal at different angles reflecting off crystallites at angles to the surface normal, which then focus back to the same point on the camera. In the present two-dimensional geometry, the crystal and camera are lines along the dispersion axis, with the crystal having a finite thickness.
@@
For simplicity and practicality, we consider a spectrometer in von Hámos geometry v. Hámos (1932) (see Fig. 1). In this geometry, a cylindrically bent crystal is used to focus the dispersed x-rays on to a single dispersion axis on the camera. By collecting all the scattered photons on to a single line, it is easier to observe events above the camera noise. As the bending process severely degrades the resolution of perfect crystals anyway, mosaic crystals are often used in von Hámos spectrometers to further increase the efficiency of the detector by taking advantage of their high reflectivity, and because the geometry allows one to utilise mosaic focusing Preston et al. (2020), as shown in Fig. 1. Due to their very high collection efficiency, mosaic crystal von Hámos spectrometers are widely-employed at FEL and laser facilities; therefore, the choice of von Hámos geometry here is relevant for typical applications of mosaic crystals, though we expect the results in this work can be generalised to alternative geometries. In von Hámos geometry, the distance from the source to a point on the camera $X_{\rm cam}$ can be readily converted to a photon energy using the dispersion equation v. Hámos (1932):
@@
where $d$ is the crystal lattice spacing, $r$ is the distance of the crystal below the source-camera axis, $h$ is Planck’s constant, and $c$ is the speed of light.
For further simplicity and efficiency, we only treat the spectrometer in two-dimensions, as is shown in Fig. 1. In other words, the source is treated as a point, the camera and the crystal are treated as lines along the dispersion axis, and the crystal has a finite thickness. We note this two-dimensional setup inherently does not explicitly treat any broadening effects off the dispersion axis on the camera and along the bent crystal. Nevertheless, as will be demonstrated, this simplified treatment of the spectrometer still provides good agreement with experimental measurements.
While we focus here on the effects of the crystal broadening, some mention should be given to the effect of the photosensitive optic on the measured spectrum. As the optic has a finite size, its most immediate effect is to force the measurement to a finite spectral window. Focusing now on CCD cameras Gruner, Tate, and Eikenberry (2002) (which tend to be used at facilities), the continuous photon energy distribution is discretized into the finite-size pixels – usually, though, the pixel size is small enough that the produced spectrum will look smooth and continuous. The photons are detected by their interaction with a e.g. Si layer in the camera: as a photon passes through the Si, it produces electron-hole (eh) pairs, with the number of eh pairs produced being proportional to the energy of the photon. The charge produced is then associated to an analogue-to-digital unit (ADU). The ADU is therefore directly proportional to photon energy and can be calibrated for individual pixels for a given camera Redford et al. (2016); Mozzanica et al. (2016). By using the dispersion relation of the spectrometer, the spectrum on the camera represents number of photons detected at a specific photon energy. The pixels also subtend a finite solid angle, which can also be readily accounted for. In short, while the camera does form part of the spectrometer IF, its effects are generally well-understood.
@@ II.1 Depth Broadening
Figure 2: The differential reflectivity from depth broadening in three scenarios involving a 40 $\mu$m thick graphite crystal, considering up to $p=4$ reflections. In all cases, the circular markers indicate the energy difference on camera that a photon can be reflected to. (a) Depth broadening for a photon with different energies incident at its Bragg angle on a crystal without mosaicity. Note that the shift on the camera is only for $\Delta E>0$, and that the maximum $\Delta E$ is different for each photon energy. (b) The same as (a), but now the crystal has a mosaicity $\gamma=0.08^{\circ}$. (c) Depth broadening for a fixed photon energy on a mosaic crystal at different angles of incidence.
Depth broadening is simply the broadening that results from the photon travelling into a finite depth of the crystal before scattering. The reflection event therefore occurs at a distance slightly further away from the source than when reflecting off the surface, so the photon will be incident on the detector slightly further away. As the finite thickness of a crystal provides more opportunities for a photon to scatter off a crystal layer, thicker crystals will have a higher integrated reflectivity at the cost of resolution.
@@
$\mathcal{R}_{p}(\alpha_{\rm in},z_{\rm eff},\gamma,T_{c},E_{0})$ $=G_{p}^{Tc}(z_{\rm eff})\sigma^{2p-1}\exp({-\mu_{\rm eff}z_{\rm eff}})\,,$ (7)
where an odd ($12p-$ 1) number of reflections is required for the photon to leave the front surface of the crystal and be detected. In Eq. (7), $\alpha_{\rm in}$ is the angle of incidence on the crystal, $z_{\rm eff}$ is the effective depth a photon travels through the crystal, $\gamma$ is the full width at half-maximum (FWHM) of the rocking curve of the crystal, and $E_{0}$ is the energy of the incident photon. $G_{p}^{Tc}(z_{\rm eff})$ is a degeneracy factor that accounts for the number of paths that produce the same $z_{\rm eff}$. The reflecting power $\sigma$ of a crystal or crystallite with thickness $dT$ is given by Zachariasen (1967):
$\sigma dT=\left|\frac{e^{2}F_{H}}{m_{e}c^{2}V}\right|^{2}\lambda^{3}\frac{1+\cos^{2}(2\alpha_{B})}{2\sin(2\alpha_{B})\sin(\alpha_{B})}W(\Delta\alpha,\gamma)dT\,,$ (8)
where $F_{H}$ is the structure factor for the Miller indices $H$, $V$ is the volume of the unit cell, $\lambda$ is the wavelength of the incident photon, $\alpha_{B}$ is the Bragg angle, $W(\Delta\alpha,\gamma)$ is the rocking curve function, and $\Delta\alpha=\alpha_{B}-\alpha_{\rm in}$. Lastly, the term $\mu_{\rm eff}$ is the effective attenuation coefficient Schlesiger et al. (2017):
$\mu_{\rm eff}=\mu_{0}\left(\frac{1}{\sin(\alpha_{\rm in})}+\frac{1}{\sin(\alpha_{\rm diff})}\right)+2\sigma\,,$ (9)
where $\mu_{0}$ is the attenuation coefficient of a photon with energy $E_{0}$ into the material, and $\alpha_{\rm diff}=2\alpha_{B}-\alpha_{\rm in}$ is the outgoing angle of the diffracted ray.
The sum of the differential reflectivity in Eq. (7) over the reflection orders $p$, $\mathcal{R}(\alpha_{\rm in},z_{\rm eff},\gamma,T_{c},E_{0})$, provides the probability a photon with an angle of incidence $\alpha_{\rm in}$ will travel an effective depth $z_{\rm eff}$ into the crystal. As these parameters also determine the path along which the photon leaves the crystal, it also provides the photon distribution on the camera.
@@ II.2 Mosaic Broadening
Figure 3: A break down of the effects contributing to mosaic broadening, just for reflections from the surface of a 80 mm long HAPG crystal. Left: A plot of Eq. (10) in von Hámos geometry for different photon energies incident on the crystal. The mosaic broadening is towards higher energies on the camera than the photon energy, and there is overlap in the position on the detector that the crystal will reflect to. The vertical dotted lines indicate the edges of the crystal. Right: The mosaic function as it appears on the camera. The drops in the reflectivity are when the edges of the crystal are reached. Note that in both plots if these energies extend beyond the camera range then there will be a hard cut as the photon can no longer be detected.
@@
As is usually the case, the higher integrated reflectivity of a mosaic crystal comes at the cost of a worse resolution as the angular distribution of the crystallites also broadens the measured spectrum – while the Bragg angle is satisfied across the crystal, the point the photon arrives on camera is still shifted from its nominal position. The mosaic broadening function has previously been considered to resemble the mosaic distribution function Preston et al. (2020). However, in von Hámos geometry , the mosaicity actually only results in broadening towards higher energies, which can be demonstrated with trigonometry: in von Hámos geometry, a scattered photon will hit the camera at a distance
$X_{\rm cam}=r\left[\cot(\alpha_{\rm in})+\cot(\alpha_{\rm diff})\right]\,,$ (10)
@@
away from the source. Hence, $X\geq X_{0}=2r\cot(\alpha_{B})$, so the crystal mosaicity will broaden a monochromatic beam only towards higher energies, which is shown in Fig. 3. This equation also implies that there are cut-offs in the spectrum when the edges of the crystal are reached as this function is not one-to-one with respect to the angle of incidence. This effect is also plotted in Fig. 3, which shows this effect in HAPG across a realistic camera range ($E_{\rm cam}\sim 7800-8700$ eV). Evidently, there are quite pronounced differences in the reflectivity across the camera range due to these cut offs.
@@
What is particularly noteworthy in Fig. 3 is that the energy broadening is extremely extended due to the Lorentzian nature of the HAPG crystallite distribution. While the distribution does cover a range of orders of magnitude, for sufficiently bright features such large deviation would be detectable: as shall be demonstrated later, the quasi-elastic signal from a monochromated beam can still be meaningfully detected out to energies $>150$ eV away. So while the most apparent broadening from a mosaic crystal may appear relatively narrow (e.g. $\sim 2.5$ eV in HAPG Preston et al. (2020)), the tails of the spectrum are still important to consider as they do not decay quickly.
@@
Another consequence of mosaic broadening is that it effects the shape of the depth broadening as well, in particular enhancing the prominence of the higher order reflections relative to the first order. An example is shown in the comparison of Fig. 2 (a) and (b), where the overall shape of the depth broadening is dramatically changed with a finite mosaicity. Indeed, the position of the first order cut-off depends on the relative thickness and mosaicity of the crystal Schlesiger et al. (2017). Also shown in Fig. 2 (c) is the depth broadening when changing the angle of incidence away from the nominal Bragg angle. Due to the mosaicity, the differential reflectivity is still finite though reduced because of the crystallite distribution function.
@@
Like the depth broadening, the mosaicity also depends on the incident photon energy though it is a quite weak dependence Zastrau et al. (2013); Gerlach et al. (2015). Overall, it is apparent that the mosaic broadening has quite a strong dependence on the spectrometer geometry, specifically the relative positions of the source, crystal, and camera. It is also clear that the mosaic and depth broadening only contribute to higher energies, meaning that the final crystal instrument function must be asymmetric towards higher photon energies.

### v2

@@ Abstract
Mosaic crystals, with their high integrated reflectivities, are widely-employed in spectrometers used to diagnose high energy density systems. X-ray Thomson scattering (XRTS) has emerged as a powerful diagnostic tool of these systems, providing in principle direct access to important properties such as the temperature via detailed balance. However, the measured XRTS spectrum is broadened by the spectrometer instrument function (IF), and without careful consideration of the IF one risks misdiagnosing system conditions. Here, we consider in detail the IF of 40 $\mu$m and 100 $\mu$m mosaic HAPG crystals, and how the broadening varies across the spectrometer in an energy range of 6.7-8.6 keV. Notably, we find a strong asymmetry in the shape of the IF towards higher energies. As an example, we consider the effect of the asymmetry in the IF on the temperature inferred via XRTS for simulated 80 eV CH plasmas, and find that the temperature can be overestimated if an approximate symmetric IF is used. We therefore expect a detailed consideration of the full IF will have an important impact on system properties inferred via XRTS in both forward modelling and model-free approaches.
@@ I Introduction
where ${\bm{q}}$ is the scattering vector, $E$ is the energy of the scattered photon measured on the spectrometer, $E_{0}$ is the energy of the incident photon, and $\hbar\omega=E_{0}-E$ is the energy loss. The SIF contains all the effects that broaden the measured XRTS spectrum, specifically the spectrum of the incident probe beam (the source) and the instrument function of the spectrometer. The temperature $k_{B}T=\beta^{-1}$ of a system in thermal equilibrium, for example, may in principle be extracted using the detailed balance relation of the DSF:
@@
An x-ray spectrometer typically consists of two components: a dispersive crystal and a photosensitive detection method (diode, scintillator, detector, etc). As the light emanating from a target is polychromatic, in order to measure the spectrum the different photon energies need to be separated from each other. This is often done using a crystal with an appropriate lattice spacing to disperse the x-rays in angle, which are then detected spatially offset in the detection plane. The spatial position of a photon on the detector can be related to its energy via Bragg’s law and the geometry of the spectrometer v. Hámos (1932); Zachariasen (1967).
Mosaic crystals are a popular choice for the dispersive crystal. While perfect crystals consist of uniform layers of atoms, mosaic crystals consist of small perfect crystallites with normal vectors randomly distributed about the surface normal Zachariasen (1967); Freund (1988); Legall et al. (2006a); Gerlach et al. (2015). As a result, in contrast to perfect crystals where the Bragg condition will only be satisfied in a certain region in the crystal depending on the rocking curve, on a mosaic crystal it can be satisfied anywhere on its surface. The solid angle coverage of a mosaic crystal is therefore substantially higher than that of a perfect crystal, resulting in a much higher integrated reflectivity Legall et al. (2006b, 2009). In the diagnostics of WDM systems in the laboratory, where these systems are often transient, this higher reflectivity has led to the widespread adoption of mosaic crystals as they collect more photons in a given shot. But this increased reflectivity comes at the cost of spectral resolution as the distribution of the crystallites also broadens the measured spectrum. For example, in widely-used Highly Annealed Pyrolytic Graphite (HAPG) crystals, the distribution of crystallites is approximately Lorentzian Zastrau et al. (2013); Gerlach et al. (2015), and therefore the crystal broadening of spectra has often been treated as a convolution of the source spectrum with a Voigt profile Preston et al. (2020); MacDonald et al. (2021).
However, recent ray tracing simulations Zastrau et al. (2012); Šmíd, Pan, and Falk (2021) and experimental data Pak et al. (2004); Döppner et al. (2008); Voigt et al. (2021) indicate that the actual broadening by mosaic crystals is very asymmetric towards higher energies. An asymmetric SIF would have consequences on conditions derived from XRTS measurements if an approximate SIF that is symmetric – or one without sufficient asymmetry – were used in its place. A clear example would be that the temperature extracted via detailed balance would be incorrect as the upshifted side of the DSF would be inferred to be larger to compensate for the increase in intensity by the SIF.
@@
In this paper, we present a systematic analysis of the broadening by mosaic crystals and the effect it has on plasma conditions inferred in XRTS measurements. To do so, we use Schlesiger et al.’s multi-reflection mosaic crystal model Schlesiger et al. (2017) to calculate the differential reflectivity along paths a photon can take from the source to the detector. The model is found to have good agreement with experimental measurements taken at the European XFEL on ambient Al and polypropylene. Notably, we find the mosaic broadening and depth broadening result in highly asymmetric broadening towards the higher energies. The instrument function is also found to be extremely extended, with substantial intensity even more than 100 eV away from the central photon energy. Furthermore, it is demonstrated that such strong asymmetry in the beam has implications on conditions inferred from XRTS via forward modelling and the model-free method if an approximate symmetric SIF is used. Finally, we investigate the validity of the approximation of the SIF as a convolution problem rather than a kernel problem.
@@ II Broadening by Mosaic Crystals
Figure 1: A two-dimensional schematic of a von Hámos detector with a mosaic crystal, with the paths of two different photon energies, $E_{1}$ and $E_{2}>E_{1}$, to the detector also shown. In full von Hámos geometry, the crystal is cylindrically bent with a radius of curvature $r$, while $X_{\rm cam}$ is the distance from the source to a point on the detector. Mosaic focusing in this geometry is represented by photons incident on the crystal at different angles reflecting off crystallites at angles to the surface normal, which then focus back to the same point on the detector. In the present two-dimensional geometry, the crystal and detector are lines along the dispersion axis, with the crystal having a finite thickness.
@@
For simplicity and practicality, we consider a spectrometer in von Hámos geometry v. Hámos (1932) (see Fig. 1). In this geometry, a cylindrically bent crystal is used to focus the dispersed x-rays on to a single dispersion axis on the detector. By collecting all the scattered photons on to a single line, it is easier to observe events above the detector noise. As the bending process severely degrades the resolution of perfect crystals anyway, mosaic crystals are often used in von Hámos spectrometers to further increase the efficiency of the detector by taking advantage of their high reflectivity, and because the geometry allows one to utilise mosaic focusing Preston et al. (2020), as shown in Fig. 1. Due to their very high collection efficiency, mosaic crystal von Hámos spectrometers are widely-employed at FEL and laser facilities; therefore, the choice of von Hámos geometry here is relevant for typical applications of mosaic crystals, though we expect the results in this work can be generalised to alternative geometries. In von Hámos geometry, the distance from the source to a point on the detector $X_{\rm cam}$ can be readily converted to a photon energy using the dispersion equation v. Hámos (1932):
@@
where $d$ is the crystal lattice spacing, $r$ is the distance of the crystal below the source-detector axis, $h$ is Planck’s constant, and $c$ is the speed of light.
For further simplicity and efficiency, we only treat the spectrometer in two-dimensions, as is shown in Fig. 1. In other words, the source is treated as a point, the detector and the crystal are treated as lines along the dispersion axis, and the crystal has a finite thickness. We note this two-dimensional setup inherently does not explicitly treat any broadening effects off the dispersion axis on the detector and along the bent crystal. Nevertheless, as will be demonstrated, this simplified treatment of the spectrometer still provides good agreement with experimental measurements.
While we focus here on the effects of the crystal broadening, some mention should be given to the effect of the detection method on the measured spectrum. As an area detector has a finite size, its most immediate effect is to force the measurement to a finite spectral window. The continuous photon energy distribution is discretized into the finite-size pixels – usually, though, the pixel size is small enough that the produced spectrum will look smooth and continuous. By using the dispersion relation of the spectrometer, the spectrum on the detector represents number of photons detected at a specific photon energy. The pixels also subtend a finite solid angle, which can also be readily accounted for, though typically the broadening introduced by this is much less than that intrinsically from bent crystals. In short, while the detector does form part of the spectrometer IF, its effects are generally well-understood.
@@ II.1 Depth Broadening
Figure 2: The differential reflectivity from depth broadening in three scenarios involving a 40 $\mu$m thick graphite crystal, considering up to $p=4$ reflections. Note the marked increase in reflectivity when mosaicity is taken into account. In all cases, the circular markers indicate the energy difference on detector that a photon can be reflected to. (a) Depth broadening for a photon with different energies incident at its Bragg angle on a crystal without mosaicity. Note that the shift on the detector is only for $\Delta E>0$, and that the maximum $\Delta E$ is different for each photon energy. (b) The same as (a), but now the crystal has a mosaicity $\gamma=0.08^{\circ}$. (c) Depth broadening for a fixed photon energy on a mosaic crystal at different angles of incidence.
Depth broadening is simply the broadening that results from the photon travelling into a finite depth of the crystal before scattering. The reflection event therefore occurs at a distance slightly further away from the source than when reflecting off the surface, so the photon will be incident on the detector slightly further away from the source. From Eq. (6), the depth broadening is therefore towards higher energies. As the finite thickness of a crystal provides more opportunities for a photon to scatter off a crystal layer, thicker crystals will have a higher integrated reflectivity at the cost of resolution.
@@
$\mathcal{R}_{p}(\alpha_{\rm in},z_{\rm eff},\gamma_{C},T_{c},E_{0})$ $=G_{p}^{Tc}(z_{\rm eff})\sigma^{2p-1}\exp({-\mu_{\rm eff}z_{\rm eff}})\,,$ (7)
where an odd ($12p-$ 1) number of reflections is required for the photon to leave the front surface of the crystal and be detected. In Eq. (7), $\alpha_{\rm in}$ is the angle of incidence of the photon on the crystal, $z_{\rm eff}$ is the effective depth a photon travels through the crystal, $T_{c}$ is the thickness of the crystal, $\gamma_{C}$ is the full width at half-maximum (FWHM) of the rocking curve of the crystal, and $E_{0}$ is the energy of the incident photon. $G_{p}^{Tc}(z_{\rm eff})$ is a degeneracy factor that accounts for the number of paths that produce the same $z_{\rm eff}$. The reflecting power $\sigma$ of a crystal or crystallite with thickness $dT$ is given by Zachariasen (1967):
$\sigma dT=\left|\frac{e^{2}F_{H}}{m_{e}c^{2}V}\right|^{2}\lambda^{3}\frac{1+\cos^{2}(2\alpha_{B})}{2\sin(2\alpha_{B})\sin(\alpha_{B})}W(\Delta\alpha,\gamma_{C})dT\,,$ (8)
where $F_{H}$ is the structure factor for the Miller indices $H$, $V$ is the volume of the unit cell, $\lambda$ is the wavelength of the incident photon, $\alpha_{B}$ is the Bragg angle for the photon energy $E_{0}$, $W(\Delta\alpha,\gamma_{C})$ is the rocking curve function, and $\Delta\alpha=\alpha_{B}-\alpha_{\rm in}$. Lastly, the term $\mu_{\rm eff}$ is the effective attenuation coefficient Schlesiger et al. (2017):
$\mu_{\rm eff}=\mu_{0}\left(\frac{1}{\sin(\alpha_{\rm in})}+\frac{1}{\sin(\alpha_{\rm out})}\right)+2\sigma\,,$ (9)
where $\mu_{0}$ is the attenuation coefficient of a photon with energy $E_{0}$ into the material, and $\alpha_{\rm out}=2\alpha_{B}-\alpha_{\rm in}$ is the angle of the outgoing diffracted ray.
The sum of the differential reflectivity in Eq. (7) over the reflection orders $p$, $\mathcal{R}(\alpha_{\rm in},z_{\rm eff},\gamma,T_{c},E_{0})$, provides the probability a photon with an angle of incidence $\alpha_{\rm in}$ will travel an effective depth $z_{\rm eff}$ into the crystal. As these parameters also determine the path along which the photon leaves the crystal, it also provides the photon distribution at the detection plane.
@@ II.2 Mosaic Broadening
Figure 3: A break down of the effects contributing to mosaic broadening, just for reflections from the surface of a 80 mm long HAPG crystal. Left: A plot of Eq. (10) in von Hámos geometry for different incident photon energies $E_{0}$, showing the energy shift on the detector versus the shift in the angle of incidence on the crystal from the photon’s Bragg angle. The mosaic broadening is towards higher energies on the detector than the photon energy, and there is overlap in the position on the detector that the crystal will reflect to. The vertical dotted lines indicate the edges of the crystal. Right: The mosaic function as it appears on the detector. The drops in the reflectivity are when the edges of the crystal are reached.
@@
As is usually the case, the higher integrated reflectivity of a mosaic crystal comes at the cost of a worse resolution as the angular distribution of the crystallites also broadens the measured spectrum – while the Bragg angle is satisfied across the crystal, the point the photon arrives on the detector is still shifted from its nominal position. The mosaic broadening function has previously been considered to resemble the mosaic distribution function Preston et al. (2020). However, in von Hámos geometry, the mosaicity actually only results in broadening towards higher energies, which can be demonstrated trigonometrically. In von Hámos geometry, a scattered photon will hit the detector at a distance:
$X_{\rm cam}$ $=r\left[\cot(\alpha_{\rm in})+\cot(\alpha_{\rm out})\right]$ (10)
$=r\left[\cot(\alpha_{B}-\Delta\alpha)+\cot(\alpha_{B}+\Delta\alpha)\right]\,,$
away from the source. This equation is plotted in the left panel of Fig. 3. When $\alpha_{\rm in}=\alpha_{B}$, the nominal position of the photon on the detector is obtained, $X_{0}=2r\cot(\alpha_{B})$. Generally speaking, $\Delta\alpha$ will be very small across a crystal, so Eq. (10) can be expanded as a Taylor series:
$X_{\rm cam}$ $=r\left[2\cot(\alpha_{B})+2\cot(\alpha_{B})\csc^{2}(\alpha_{B})\Delta\alpha^{2}+\mathcal{O}(\Delta\alpha^{4})\right]$ (11)
$=X_{0}+X_{0}\csc^{2}(\alpha_{B})\Delta\alpha^{2}+\mathcal{O}(\Delta\alpha^{4})\,.$
From this quadratic, it is clear $X_{\rm cam}\geq X_{0}$, for all $\alpha_{B}$ and $\Delta\alpha$, which is what is observed in the left panel of Fig. 3.
@@
The relationship in Eq. (10) also implies that there are cut-offs in the spectrum when the edges of the crystal are reached as this function is not one-to-one with respect to the angle of incidence. This effect is plotted in the right panel of Fig. 3, which shows this effect in HAPG across a realistic detector range ($E_{\rm cam}\sim 7800-8700$ eV). Evidently, there are quite pronounced differences in the reflectivity across the detector range due to these cut offs.
@@
What is particularly noteworthy in Fig. 3 is that the energy broadening is extremely extended due to the Lorentzian nature of the HAPG crystallite distribution. While the distribution does cover a range of orders of magnitude, for sufficiently bright features such a large deviation would be detectable: as shall be demonstrated later, the quasi-elastic signal from a monochromated beam can still be meaningfully detected out to energies $>150$ eV away. So while the most apparent broadening from a mosaic crystal may appear relatively narrow (e.g. $\sim 2.5$ eV in HAPG Preston et al. (2020)), the tails of the spectrum are still important to consider as they do not decay quickly.
@@
Another consequence of mosaic broadening is that it affects the shape of the depth broadening as well, in particular enhancing the prominence of the higher order reflections relative to the first order. An example is shown in the comparison of Fig. 2 (a) and (b), where the overall shape of the depth broadening is dramatically changed with a finite mosaicity. Indeed, the position of the first order cut-off depends on the relative thickness and mosaicity of the crystal Schlesiger et al. (2017). Also shown in Fig. 2 (c) is the depth broadening when changing the angle of incidence away from the nominal Bragg angle. Due to the mosaicity, the differential reflectivity is still finite though reduced because of the crystallite distribution function.
@@
Like the depth broadening, the mosaicity also depends on the incident photon energy though it is a quite weak dependence Zastrau et al. (2013); Gerlach et al. (2015). Overall, it is apparent that the mosaic broadening has quite a strong dependence on the spectrometer geometry, specifically the relative positions of the source, crystal, and detector. It is also clear that the mosaic and depth broadening only contribute to higher energies, meaning that the final crystal instrument function must be asymmetric towards higher photon energies.